[Brahms-l] Shift report 20040329 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Mon Mar 29 2004 - 16:08:07 EST
------------------------------------------------------------
Date: Mon Mar 29 16:08:05 2004, Shift: 08:00-16:00
Supervisor: ykl, with: 
Log book#15 pages 101 - 101

Shift Summary:
------------------------------------------------------------
1. A store for 4 hours and another store for three hours
and still continueing.

2. MRS at 45 degrees 6KGB, and FS at 6 degrees 1/4A.
Run 11397 has much background and FV wants to discredit.

Run 11399 through 11415 with many runs to debug DAQ
problems and pedestals.

3. T2 anode A is restored to the default voltage 1290.
Daq seems to repeat "time out" problems. See the remedy
in the e-log.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Mar 29 16:08:31 2004

This archive was generated by hypermail 2.1.8 : Mon Mar 29 2004 - 16:08:54 EST